A rectangle has an area of 14 square centimeters. It's length is five cm greater than its width. Find the dimensions.
Found 2 solutions by Alan3354, jorel1380:
Answer by Alan3354(69443)

A rectangle has an area of 14 square centimeters. It's [sic] length is five cm greater than its width. Find the dimensions.
-----------------
Find a pair of factors of 14 that differ by 5.
----
it's = it is

Answer by jorel1380(3719)

Let n be the width of the rectangle, in cm. Then the length would be w+5. So:
w(w+5)=14
w²+5w-14=0
(w+7)(w-2)=0
w=2 or -7
We get a positive value of 2cm. for the width, and 7cm. for the length. ☺☺☺☺